Hello, I was hoping you could help me, I have tried and tried but I can't work out how to isolate the x to find its value, please help, and thank you in advance!
The sides of a right angled trienagle are x, (x+2) and (2x-2). The hypotenuse is the length (2x-2). Find the actual dimensions of the triangle in cm.
Thank you!

Answer by FrankM(1040)

x^2 + (x+2)^2 = (2x-2)^2 is the set up for this equation.
x^2+x^2+4x+4= 4x^2-8x+4 expanded both sides
2x^2+4x+4=4x^2-8x+4
-2x^2+12x=0
2x^2-12x=0
x^2-6x=0
x=0 or x=6
6, 8, 10 are the sides